# Entanglement patterns of quantum chaotic Hamiltonians with a scalar U(1) charge

### v1

@@ Abstract
Our current understanding of quantum chaos hinges on the random matrix behavior (RMT) of typical states in quantum many-body systems, particularly eigenstates and their energy level statistics. Although RMT has been remarkably successful in describing ‘coarse’ features of quantum states in chaotic regimes, it fails to capture their ‘finer’ features, particularly those arising from spatial locality and symmetries. Here, we show that we can accurately describe the behavior of eigenstate ensembles in physical systems by using RMT ensembles with constraints that capture the key features of the physical system. We demonstrate our approach on local spin Hamiltonians with a scalar U(1) charge. By constructing constrained RMT ensembles that account for two local scalar charges playing the role of energy and magnetization, we describe the patterns of entanglement of mid-spectrum eigenstates at all lengthscales and beyond their average behavior, analytically and numerically. When defining the correspondence between quantum chaos and RMT, our work clarifies that RMT ensembles must be constrained to account for all the features of the underlying Hamiltonian, particularly spatial locality and symmetries.
@@
Introduction.—Understanding how statistical mechanics emerges in isolated quantum systems has been a long-standing challenge [1, 2, 3, 4]. The fundamental difficulty in addressing this challenge has been defining a rigorous notion of chaos in quantum regimes governed by linear unitary evolution. However, valuable insights into our understanding of quantum chaos come from random matrix theory (RMT), which has been remarkably successful in describing ‘coarse’ features of quantum states in chaotic regimes. In particular, the widely accepted expectation is that the eigenstates and eigenvalues of chaotic systems exhibit universal behavior described by RMT. This expectation applies to eigenspectrum properties, such as the level spacing statistics [5, 6, 7] or the spectral form factor [8, 9, 10], and to eigenstate properties, such as the thermal-like behavior of local observables [11, 12, 13, 14] and volume-law behavior of the entanglement entropy (EE) [15, 16, 17, 18, 19].
@@
Figure 1: (a) Histograms of half-system entanglement entropy (EE) of mid-spectrum eigenstates for Hamiltonian systems with a scalar U(1) charge (left, red) and for a Hamiltonian system without additional symmetries (right, blue). The lines indicate the EE distribution of pure random states without constraints (dotted), random states with a U(1) constraint (dashed-dotted), and random states with two U(1) constraints (dashed). (b) Schematics showing the coarse-graining procedure used to generate RMT ensembles that capture the entanglement patterns of a spin chain Hamiltonian of spin-$1/2$ qubits, shown for $L=16$. In particular, we consider a chain of $L^{\prime}=L/\lambda$ qudits with local Hilbert space dimension $d^{\prime}=d^{\lambda}$, with $\lambda=2$, and consider random states within a fixed symmetry sector for each U(1) charge. Parameters used in the histograms: $L=16$, $J=1$, $\Delta=J/2$, $J_{2}=0.6$, $h_{x}=0$ (Ham×$\times$×U(1)), and $h_{x}=0.75$, $J_{2}=0$ (Ham), see Eqs.(Entanglement patterns of quantum chaotic Hamiltonians with a scalar U(1) charge) and (8).
@@
While RMT captures ‘coarse’ features of quantum states in chaotic regimes, it fails to describe their ‘fine’ features. In particular, eigenstates of physical Hamiltonians encode correlations, whereas RMT ensembles do not. In fact, the correlations encoded in a single eigenstate is sufficient to reconstruct the entire Hamiltonian, so long as the Hamiltonian is local [20, 15]. Many recent works have aimed to capture structure in quantum states that goes beyond RMT. For example, Refs. [21, 22, 23, 24, 25, 26, 27] showed that matrix elements of local operators evaluated in the eigenbasis of the Hamiltonian remain correlated up to a particular energy scale related to the so-called Thouless time [8, 28, 29, 9, 10]. These works clarified that correlations are related to an emergent “light cone” in the growth of out-of-time ordered commutators in spatially local systems [25, 26, 27]. Other works have identified systematic deviations between the EE of mid-spectrum Hamiltonian eigenstates and pure random states [30, 31, 32, 33, 17, 34].
@@
Despite all these works, notions of quantum chaos that systematically incorporate additional structure beyond RMT are lacking. The perspective that we take in this letter is to systematically construct RMT ensembles that capture increasingly finer features of typical quantum states of physical Hamiltonians by imprinting additional structure on quantum states, such as spatial locality and symmetries. An initial step in this direction was done in our recent work [35] where we showed that the EE of mid-spectrum eigenstates of local Hamiltonian systems (in the absence of any additional structure) is not described by plain RMT but, instead, by a constrained RMT ensemble that accounts for a U(1) scalar charge [35]—energy in local systems plays the role of the scalar charge. Similar arguments were put forward by Huang [31, 32] to justify deviations from the maximal EE of eigenstates in local Hamiltonian systems.
@@
More generally, here we demonstrate that we can accurately describe the entanglement patterns of eigenstates of physical systems—including their fluctuations—by appropriately constraining the RMT distribution, focusing on Hamiltonian systems with an additional U(1) scalar charge, see Fig.1. Systems with a U(1) scalar charge arise in models of experimental interest, such as magnetic systems, and systems with particle number conservation, such as interacting Fermi and Bose gases [36, 37]. By systematically constructing RMT ensembles that account for conservation of energy and magnetization (or particle number), we accurately describe the ensemble properties of eigenstates beyond their average behavior, both analytically and numerically. Specifically, using this approach, we describe the entanglement patterns of mid-spectrum eigenstates, including their fluctuations, at all lengthscales (Figs.1 and 2).
Ensembles without any constraints.—We consider ensemble properties of eigenstates and pure random states through the lens of EE,
$S_{A}=-\text{Tr}[\rho_{A}\log(\rho_{A})],$ (1)
in the context of spin chains comprised of $L$ qudits with local Hilbert space dimension $d$. In Eq.(1), $\rho_{A}={\rm Tr}_{B}[|\Psi\rangle\langle\Psi|]$ is the reduced density matrix of the state $|\Psi\rangle$ when the system is bipartitioned into two subsystems of length $L_{A}=fL$ and $L_{B}=(1-f)L$, each of which has Hilbert space dimension $d_{A}=d^{L_{A}}$ and $d_{B}=d^{L_{B}}$. An ensemble of quantum states $\{|\Psi\rangle\}$ produces a distribution $\mathcal{P}(S_{A})$ of EE, which we aim to describe at the level of higher moments.
For example, in the absence of any structure, the distribution of EE of pure random states $|\Psi\rangle$ drawn from the full Hilbert space $\in{\cal H}$ depends only on subsystem dimensions through the parameters $(f,L)$. In particular, the average EE, $\mu=\langle S_{A}\rangle$, in the asymptotic limit ($L\gg 1$) is given by
@@
which was first conjectured by Page [38] and later proven analytically by others [39, 40, 41]. The first term in the RHS of Eq.(2) is the volume-law term which scales with subsytem size $L_{A}=fL$, whereas the second term gives rise to the ‘half-qubit’ shift correction for half-subsystems. The variance of EE for pure random states, $\sigma^{2}_{P}\approx d^{-L(1+|1-2f|)}$, is exponentially small in subsystem size, implying that the EE is typical and a single pure random state will have the Page entropy.
Ensembles constrained by one scalar charge.—Energy conservation in local Hamiltonian systems adds structure into eigenstates which is not described by plain RMT. Although there are subtle differences between energy conservation and U(1) charge conservation, namely, the energy operator is sum of $k$-local terms (typically, $k=2$ or $3$) with a continuously distributed eigenspectrum, we expect that such microscopic differences are washed away when computing the EE for large enough subsystems ($d_{A}\gg 1$). Such arguments were discussed in more detail in our work [35].
For systems with a local scalar charge and a local Hilbert space dimension of $d=2$, it is convenient to think of $0\leq M\leq L$ as an integer particle number, and each site only able to accommodate a maximum of one particle. The Hilbert space ${\mathcal{H}(M)}$ of states with fixed charge $M$ decomposes as a direct sum of tensor products, ${\cal H}(M)=\bigoplus_{M_{A}}{\cal H}_{A}(M_{A})\otimes{\cal H}_{B}(M-M_{A})$, where $M_{A}$ is within the range ${\rm max}(0,M-L_{B})\leq M_{A}\leq{\rm min}(M,L_{A})$. The Hilbert space dimension of ${\cal H}_{A}(M_{A})$ is $d_{A,M_{A}}=\binom{L_{A}}{M_{A}}$, the Hilbert space dimension of ${\cal H}_{\rm B}(M-M_{A})$ is $d_{B,M-M_{A}}=\binom{L-L_{A}}{M-M_{A}}$, and the total Hilbert space dimension is $\sum_{M_{A}}d_{A,M_{A}}d_{B,M-M_{A}}=d_{M}=\binom{L}{M}$. A random state with fixed total charge $|\Psi_{M}\rangle\in{\cal H}(M)$ can be expressed as a superposition of orthonormal basis states, $|\Psi_{M}\rangle=\sum_{M_{A}}\psi_{\alpha,\beta}^{(M_{A})}|M_{A},\alpha\rangle\otimes|M-M_{A},\beta\rangle$, with $\psi^{(M_{A})}_{\alpha,\beta}$ uncorrelated random numbers up to normalization. The index $\alpha$ ($\beta$) labels the basis states in subsystem $A$ ($B$) with a total charge $M_{A}$ ($M-M_{A}$).
@@
Figure 2: (a) Finite-size scaling of the EE distribution as a function of $L$ for fixed $f=1/2$. The data is shown relative to the maximum entropy, $\delta S_{A}=Lf\log(2)-S_{A}$. The red points indicate the average EE, and the bars indicate the standard deviation of EE. (b) Distribution of EE for mid-spectrum eigenstates as a function of $f$ for fixed $L=16$. The shaded region in panels (a) and (b) are obtained from sampling pure random states with a U(1) constraint (blue) and a U(1)$\times$×U(1) constraint (red) in coarse-grained systems of size $L^{\prime}=[4,6,8]$, see Fig.1(b). Parameters used in panels (a) and (b): (red) $J=1$, $\Delta=J/2$, $J_{2}=0.6$, $h_{x}=0$ and (blue) $J_{2}=0$, $h_{x}=0.75$. For small values of $f$, the standard deviation bars are hidden behind the data points.
@@
The first few moments of the EE distribution produced by the ensemble $|\Psi\rangle\in{\cal H}(M)$ was first computed by Bianchi and Dona [41], see details in the Supplement. In particular, the mean entanglement entropy for ‘mid-spectrum’ states (i.e., $M/L=1/2$) in the asymptotic limit is given by
$\mu_{M}(L,f)=Lf\log(d)+\frac{f+\log(1-f)}{2}-\frac{1}{2}\delta_{f,1/2}.$ (4)
@@
Interestingly, in addition to the volume-law term and the half-qubit shift, Eq.(4) also exhibits a finite shift in the mean EE entropy relative to the Page result [38]. The variance of EE scales exponentially with system size, $\sigma_{M}\sim\sqrt{L}/d^{L}$, thus a typical pure random state in ${\cal H}(M)$ will have the EE in Eq.(4). We emphasize that the difference between the typical EE of random states in ${\cal H}(M)$ and ${\cal H}$ is significant on the exponentially small scale set by $\sigma_{M}$, see dotted and dashed-dotted lines in Fig.1(a).
We note that if random vectors are real valued (i.e., GOE distributed), which describes states in systems with time-reversal symmetry, the mean EE is not affected at the level of $O(1)$ corrections, but the standard deviation increases by a factor of $\sqrt{2}$ [42, 39]. This is true both for constrained and unconstrained ensembles (see Supplement).
@@
Ensembles constrained by two commuting scalar charges.—We now consider quantum state ensembles constrained by an additional U(1) charge, which is more descriptive of typical quantum states in systems that conserve energy and magnetization. Unlike the U(1) case, there is a technical difficulty in constructing random state ensembles that have two scalar constraints if the local Hilbert space dimension is $d=2$: the symmetry operators for both charges cannot be simultaneously expressed as a sum of single site terms, thus defining a suitable basis to describe the states is challenging. To address this, we use a ‘coarse-graining’ picture in which we increase the local Hilbert space dimension as $d\rightarrow d^{\lambda}$, and decrease the system size as $L\rightarrow L/\lambda$, while keeping the total Hilbert space dimension constant. By ‘enlarging’ the local Hilbert space dimension, we can express the symmetry operators for both U(1) charges in terms of local operators only. Similarly to the energy-conserving-only case, we argue that the entanglement patterns computed for large enough subsystems are insensitive to the coarse-graining procedure—and, in fact, we find that the entanglement patterns are also insensitive when the subsystem is not so large.
@@
For systems constrained by two scalar U(1) charges and a local Hilbert space dimension of $d=2$, we use $\lambda=2$ and define the random states ensemble, $|\Psi_{M,N}\rangle=\sum_{M_{A},N_{A}}\psi_{\alpha,\beta}^{(M_{A},N_{A})}|M_{A},N_{A};\alpha\rangle\otimes\\ |M-M_{A},N-N_{A};\beta\rangle$, with local Hilbert space dimension $d^{\prime}=4$. The numbers $N$ and $M$ represent the quantum numbers of each scalar charge, and each site can accommodate one particle of each flavor, see Fig.1(b). The index $\alpha$ ($\beta$) labels basis states with $M_{A}$ and $N_{A}$ ($M-M_{A}$ and $N-N_{A}$) particles in $A$ ($B$). In this case, the EE of $\rho_{A}$ can be expressed as
@@
where each term has the same physical meaning as those in Eq.(3). When states are drawn randomly from ${\cal H}(M,N)$, the resulting distribution is computed along the same lines as in the U(1) case, and the details are discussed in the Supplement. We find that the mean EE for mid-spectrum states ($M/L=1/2$ and $N/L=1/2$) in the asymptotic limit is given by,
$\mu_{M,N}(L,f)=Lf\log(d)+f+\log(1-f)-\frac{1}{2}\delta_{f,1/2}.$ (6)
In particular, the mean EE is shifted twice the value found in Eq.(4) relative to the Page mean. Note that the mean EE does not change if we rescale $L\rightarrow L/\lambda$ and $d\rightarrow d^{\lambda}$, so long as $f$ remains constant. The variance of EE is exponentially small in systems size, $\sigma_{M,N}\sim L/d^{L}$ (see Supplement). We note that the ratio $\sigma_{M,N}/\sigma_{M}\sim L^{1/4}$; this leads to $\sigma_{M,N}$ being twice as large than $\sigma_{M}$ when $L=16$, as shown in Fig.1(a). Having described the three reference RMT ensembles, we now compare these with ensembles of eigenstates of Hamiltonian systems with and without U(1) symmetry.
Model Hamiltonian.—We consider the spin-1/2 Hamiltonian:
@@
where $(X_{\ell},Y_{\ell},Z_{\ell})$ are the Pauli matrices. The Hamiltonian (Entanglement patterns of quantum chaotic Hamiltonians with a scalar U(1) charge) has two scalar charges, energy and total magnetization $S_{z}=\sum_{\ell}Z_{\ell}$, both of which are commuting, $[H,S_{z}]=0$. The Hamiltonian (Entanglement patterns of quantum chaotic Hamiltonians with a scalar U(1) charge) also has multiple point symmetries, which we explicitly break. In particular, we use open boundary conditions to break translation symmetry, and we include boundary fields $h_{1}=Z_{1}/4$ and $h_{L}=-Z_{L}/4$ to break inversion symmetry. When $J_{2}=0$, the Hamiltonian (Entanglement patterns of quantum chaotic Hamiltonians with a scalar U(1) charge) becomes the integrable XXZ chain. A finite value of $J_{2}$, instead, breaks integrability. We first consider the parameters $J=1$, $\Delta=J/2$, and $J_{2}=0.6$, which we find to be the values in which eigenstates are most random.

### v2

@@ Abstract
Our current understanding of quantum chaos in many-body quantum systems hinges on the random matrix theory (RMT) behavior of eigenstates and their energy level statistics. Although RMT has been remarkably successful in describing ‘coarse’ features of many-body quantum Hamiltonians in chaotic regimes, such as the Wigner-Dyson level spacing statistics or the volume-law behavior of eigenstate entanglement entropy, it remains a challenge to describe their ‘finer’ features, particularly those arising from spatial locality. Here, we show that we can accurately describe the statistical behavior of eigenstate ensembles in many-body Hamiltonians by using pure random states with physical constraints that capture the essential features of the Hamiltonian, specifically spatial locality and symmetries. We demonstrate our approach on local spin Hamiltonians with a scalar U(1) charge. By constructing ensembles of constrained random states that account for two commuting scalar charges playing the role of energy and magnetization, we describe the patterns of entanglement of mid-spectrum eigenstates beyond their average volume-law behavior, including $O(1)$ corrections and fluctuations, analytically and numerically. When defining the correspondence between quantum chaotic eigenstates in many-body Hamiltonians and RMT ensembles, our work highlights the important role played by spatial locality in describing universal features beyond the volume-law behavior.
@@
Introduction.—Understanding how statistical mechanics emerges in isolated many-body quantum systems has been a long-standing challenge [1, 2, 3, 4]. The fundamental difficulty in addressing this challenge is describing how chaos emerges from unitary evolution in quantum systems. Valuable insights into this foundational question come from random matrix theory (RMT), which has been remarkably successful in describing ‘coarse’ features of many-body quantum states in chaotic regimes. In particular, the widely-accepted expectation is that the eigenstates and eigenvalues of generic quantum many-body Hamiltonians (i.e., away from fine-tuned integrable limits) exhibit universal statistics described by RMT ensembles, such as the Gaussian Orthogonal Ensemble (GOE) in the context of time-reversal invariant Hamiltonians. This expectation applies to eigenspectrum properties, such as the level spacing statistics [5, 6, 7] or the spectral form factor [8, 9, 10], and to eigenstate properties, such as the thermal-like behavior of local observables [11, 12, 13, 14] and the volume-law behavior of the entanglement entropy (EE) [15, 16, 17, 18, 19].
@@
Figure 1: (a) Histograms of half-system entanglement entropy (EE) of mid-spectrum eigenstates for Hamiltonian systems with a scalar U(1) charge (left, red) and for a Hamiltonian system without additional symmetries (right, blue). The lines indicate the EE distribution of pure random states without constraints (dotted), random states with a U(1) constraint (dashed-dotted), and random states with two U(1) constraints (dashed). (b) Schematics showing the coarse-graining procedure used to generate quantum state ensembles that capture the entanglement patterns of a spin chain Hamiltonian of spin-1/21/21 / 2 qubits, shown for $L=16$. In particular, we consider a chain of $L^{\prime}=L/\lambda$ qudits with local Hilbert space dimension $d^{\prime}=d^{\lambda}$, with $\lambda=2$, and consider random states within a fixed symmetry sector for each U(1) charge. Parameters used in the histograms: $L=16$, $J=1$, $\Delta=J/2$, $J_{2}=0.6$, $h_{x}=0$ [Ham×$\times$×U(1)], and $h_{x}=0.75$, $J_{2}=0$ (Ham), see Eqs.(Entanglement patterns of quantum chaotic Hamiltonians with a scalar U(1) charge) and (9).
@@
While RMT ensembles successfully describe ‘coarse’ features of quantum states in chaotic regimes, describing their ‘finer’ features remains a challenge. In particular, eigenstates of many-body Hamiltonians encode physical correlations that are not described by eigenstates of dense, random matrices such as GOE. In recent years, many works have aimed to capture structure in many-body quantum states that goes beyond RMT, particularly that arising from spatial locality. For example, Refs. [20, 21, 22, 23, 24, 25, 26] showed that matrix elements of local operators evaluated in the eigenbasis of the Hamiltonian remain correlated up to a particular energy scale related to the so-called Thouless time [8, 27, 28, 9, 10]. Other works have identified systematic deviations between the EE of mid-spectrum energy eigenstates and pure random states in spatially local systems [29, 30, 31, 32, 17, 33]. In fact, spatial locality is strongly imprinted in the structure of individual eigenstates, as a single eigenstate is sufficient to reconstruct the entire Hamiltonian [34, 15].
@@
Despite these works, statistical descriptions of many-body eigenstates in chaotic regimes that incorporate spatial locality as the fundamental building block still remain to be developed. The perspective that we take in this letter is to construct ensembles of pure random states that capture increasingly finer features of typical quantum states of physical Hamiltonians by imprinting symmetries on top of spatial locality. A first step in this direction was taken in our recent work [35], where we showed that the entanglement entropy (EE) statistics of mid-spectrum eigenstates in local many-body Hamiltonians is not described by the GOE ensemble but, instead, by a constrained ensemble that incorporates a U(1) scalar charge—energy in local systems plays the role of the scalar charge [35]. Similar arguments were used by Huang [30, 31] to justify deviations from the maximal EE of eigenstates in local Hamiltonian systems. Using this mapping between energy eigenstates and ensembles constrained by U(1) scalar charge, it is possible to derive analytic expressions for the asymptotic behavior of the EE entropy distributions, including $O(1)$ corrections and fluctuations, as calculations can be more easily done in the U(1) case.
@@
Here we show that we can accurately describe the entanglement patterns of many-body eigenstates in local systems—including $O(1)$ corrections and fluctuations—in more general settings. We focus specifically on the simplest case of Hamiltonian systems with an additional U(1) scalar charge that commutes with energy, see Fig. 1. Systems with a U(1) scalar charge arise in many models of experimental interest, such as magnetic systems, and systems with particle number conservation, such as interacting Fermi and Bose gases [36, 37]. By constructing random state ensembles that account for the conservation of energy and magnetization on top of spatial locality, we accurately describe the ensemble properties of eigenstates beyond their average volume-law behavior, both analytically and numerically (Figs. 1 and 2). Importantly, our results suggest the existence of universality in the $O(1)$ corrections and fluctuations of the EE distribution which is governed by the number of conserved charges, as discussed below.
Overview of the approach.—Our goal is to describe the statistical properties of midspectrum eigenstates in local systems through the lens of EE,
$S_{A}=-\text{Tr}[\rho_{A}\log(\rho_{A})],\quad\rho_{A}={\rm Tr}_{B}[|\Psi\rangle\langle\Psi|].$ (1)
In Eq.(1), $\rho_{A}$ is the reduced density matrix of the state $|\Psi\rangle$ when the system is bipartitioned into two subsystems of length $L_{A}=fL$ and $L_{B}=(1-f)L$, each of which has Hilbert space dimension $d_{A}=d^{L_{A}}$ and $d_{B}=d^{L_{B}}$ ($d$ being the local Hilbert space dimension). Without loss of generality, we assume $f\leq 1/2$. The states $|\Psi\rangle$ are drawn from a small microcanonical window of size $\epsilon\ll 1$ near the middle of the spectrum for each conserved quantity ${Q_{\alpha}}=\sum_{i=1}^{L}{q}_{\alpha,i}$, which we assume to be local (i.e., $q_{\alpha,i}$ acts on a few neighboring degrees of freedom) and commuting with each other.
Since we are specifically interested in the asymptotic behavior of the subsystem EE (i.e., $d_{A},d_{B}\gg 1$), the essence of our approach is based on the observation that the distribution of EE remains largely unaffected by the specific details of the local charges $Q_{\alpha}$ as the system scales up in size. This holds as long as the oprators ${q}_{\alpha,i}$ remain local, acting only on a few degrees of freedom. As such, we can adopt the simplest possible form for the local charges $q_{\alpha,i}$—specifically, single-site operators—that preserve all Hamiltonian symmetries when computing the asymptotic behavior of EE, including $O(1)$ corrections and fluctuations. For the remaining of the letter, we first describe the main analytical results as the number of conservation laws increases, followed by numerical results on Hamiltonian systems.
[…]
@@
which was first conjectured by Page [38] and later proven analytically by others [39, 40, 41]. The first term in the RHS of Eq.(2) is the volume-law term which scales with subsytem size $L_{A}=fL$, whereas the second term gives rise to the ‘half-qubit’ shift correction for half-subsystems. Higher statistical moments can be computed using a variety of methods[39, 40, 41, 42, 43]. In particular, the variance of EE for pure random states, $\sigma^{2}_{P}\approx d^{-L(1+|1-2f|)}$, is exponentially small in subsystem size, implying that the EE is typical and a single pure random state will have the Page entropy.
Ensembles constrained by one scalar charge.— For systems with a local scalar charge and a local Hilbert space dimension of $d=2$, it is convenient to think of $0\leq M\leq L$ as an integer particle number, and each site only able to accommodate a maximum of one particle. The Hilbert space ${\mathcal{H}(M)}$ of states with fixed charge $M$ decomposes as a direct sum of tensor products, ${\cal H}(M)=\bigoplus_{M_{A}}{\cal H}_{A}(M_{A})\otimes{\cal H}_{B}(M-M_{A})$, where $M_{A}$ is within the range ${\rm max}(0,M-L_{B})\leq M_{A}\leq{\rm min}(M,L_{A})$. The Hilbert space dimension of ${\cal H}_{A}(M_{A})$ is $d_{A,M_{A}}=\binom{L_{A}}{M_{A}}$, the Hilbert space dimension of ${\cal H}_{\rm B}(M-M_{A})$ is $d_{B,M-M_{A}}=\binom{L-L_{A}}{M-M_{A}}$, and the total Hilbert space dimension is $\sum_{M_{A}}d_{A,M_{A}}d_{B,M-M_{A}}=d_{M}=\binom{L}{M}$. A random state with fixed total charge $|\Psi_{M}\rangle\in{\cal H}(M)$ can be expressed as a superposition of orthonormal basis states, $|\Psi_{M}\rangle=\sum_{M_{A}}\psi_{\alpha,\beta}^{(M_{A})}|M_{A},\alpha\rangle\otimes|M-M_{A},\beta\rangle$, with $\psi^{(M_{A})}_{\alpha,\beta}$ uncorrelated random numbers up to normalization. The index $\alpha$ ($\beta$) labels the basis states in subsystem $A$ ($B$) with a total charge $M_{A}$ ($M-M_{A}$).
@@
Figure 2: (a) Finite-size scaling of the EE distribution as a function of $L$ for fixed $f=1/2$. The data is shown relative to the maximum entropy, $\delta S_{A}=Lf\log(2)-S_{A}$. The red points indicate the average EE, and the bars indicate the standard deviation of EE. (b) Distribution of EE for mid-spectrum eigenstates as a function of $f$ for fixed $L=16$. The shaded areas indicate the regions limited by $S_{A}=\mu_{M}\pm\sigma_{M}$ (blue) and $S_{A}=\mu_{M,N}\pm\sigma_{M,N}$ (red) for the ensembles with one and two scalar constraints, respectively, in coarse-grained systems of size $L^{\prime}=[4,6,8]$ [in panel (b), the shaded regions are narrow]. Parameters used in panels (a) and (b): (red) $J=1$, $\Delta=J/2$, $J_{2}=0.6$, $h_{x}=0$ and (blue) $J_{2}=0$, $h_{x}=0.75$. For small values of $f$, the standard deviation bars are hidden behind the data points.
@@
The first few moments of the EE distribution produced by the ensemble $|\Psi\rangle\in{\cal H}(M)$ was first computed by Bianchi and Dona [41], see details in the Supplement. In particular, the mean entanglement entropy for ‘mid-spectrum’ states ($M/L=1/2$) in the asymptotic limit[19] is given by
$\mu_{M}(L,f)=Lf\log 2+\frac{f+\log(1-f)}{2}-\frac{1}{2}\delta_{f,1/2}.$ (4)
@@
Interestingly, in addition to the volume-law term and the half-qubit shift, Eq.(4) also exhibits a finite shift in the mean EE entropy[17] relative to the Page result [38]. The variance of EE scales exponentially with system size, $\sigma_{M}^{2}\sim\sqrt{L}/d^{L}$, thus a typical pure random state in ${\cal H}(M)$ will have the EE in Eq.(4). We emphasize that the difference between the typical EE of random states in ${\cal H}(M)$ and ${\cal H}$ is large on the exponentially small scale set by $\sigma_{M}$, see dotted and dashed-dotted lines in Fig.1(a). We also note that if random vectors are real valued (i.e., GOE distributed), the mean EE is not affected at the level of $O(1)$ corrections, but the standard deviation increases by a factor of $\sqrt{2}$ [44, 39]. This is true both for constrained and unconstrained ensembles (see Supplement).
@@
Ensembles constrained by two commuting scalar charges.—We now consider quantum state ensembles constrained by an additional U(1) charge, which is more descriptive of typical quantum states in systems that conserve energy and magnetization. Unlike the U(1) case, there is a technical difficulty in constructing random state ensembles that have two scalar constraints if the local Hilbert space dimension is $d=2$: the symmetry operators for both charges cannot be simultaneously expressed as a sum of single site terms, thus defining a suitable basis to compute the asymptotic behavior of the EE is challenging. To address this challenge, we use a ‘coarse-graining’ procedure in which we increase the local Hilbert space dimension as $d\rightarrow d^{\lambda}$, and decrease the system size as $L\rightarrow L/\lambda$, while keeping the total Hilbert space dimension constant. By ‘enlarging’ the local Hilbert space dimension, we can express the symmetry operators for both U(1) charges in terms of local operators only. Similarly to the energy-conserving-only case, we argue that the entanglement patterns computed for large enough subsystems are insensitive to the coarse-graining procedure.
For systems constrained by two scalar U(1) charges and a local Hilbert space dimension of $d=2$, we use $\lambda=2$ and define the random states ensemble,
$|\Psi_{M,N}\rangle=\sum_{M_{A},N_{A}}\psi_{\alpha,\beta}^{(M_{A},N_{A})}|M_{A},N_{A};\alpha\rangle\otimes\\ |M_{B},N_{B};\beta\rangle,$ (5)
@@
with local Hilbert space dimension $d^{\prime}=4$, $M_{B}=M-M_{A}$, and $N_{B}=N-N_{A}$. The numbers $N$ and $M$ represent the quantum numbers of each scalar charge, and each site can accommodate one particle of each flavor, see Fig.1(b). The index $\alpha$ ($\beta$) labels basis states with $M_{A}$ and $N_{A}$ ($M_{B}$ and $N_{B}$) particles in $A$ ($B$). In this case, the EE of $\rho_{A}$ can be expressed as
@@
where each term has the same physical meaning as those in Eq.(3). When states are drawn randomly from ${\cal H}(M,N)$, the resulting distribution is computed using the same approach as in the U(1) case, and the details are discussed in the Supplement. We find that the mean EE of mid-spectrum states ($M/L=1/2$ and $N/L=1/2$) in the asymptotic limit is given by,
$\mu_{M,N}(L,f)=Lf\log 2+f+\log(1-f)-\frac{1}{2}\delta_{f,1/2}.$ (7)
In particular, the mean EE is shifted twice the value found in Eq.(4) relative to the Page mean. Note that the mean EE does not change if we rescale $L\rightarrow L/\lambda$ and $d\rightarrow d^{\lambda}$, so long as $f$ remains constant. The variance of EE is exponentially small in systems size, $\sigma_{M,N}^{2}\sim L/d^{L}$ (see Supplement). We note that the ratio $\sigma_{M,N}/\sigma_{M}$ is $\sim L^{1/4}$; this leads to $\sigma_{M,N}$ being twice as large as $\sigma_{M}$ when $L=16$, as shown in Fig.1(a). Having described the three reference ensembles, we now compare these with the eigenstate ensemble of local Hamiltonians with and without U(1) symmetry.
Eigenstate ensembles.—We consider the spin-1/2 Hamiltonian:
@@
where $(X_{\ell},Y_{\ell},Z_{\ell})$ are the Pauli matrices. The Hamiltonian (Entanglement patterns of quantum chaotic Hamiltonians with a scalar U(1) charge) has two scalar charges, energy and total magnetization, $S_{z}=\sum_{\ell}Z_{\ell}$, both of which are commuting, $[H,S_{z}]=0$. The Hamiltonian (Entanglement patterns of quantum chaotic Hamiltonians with a scalar U(1) charge) also has multiple point symmetries, which we explicitly break. In particular, we use open boundary conditions to break translation symmetry, and we include the boundary term $H_{\rm b}=h_{b}(Z_{1}-Z_{L})$ to break inversion symmetry. When $J_{2}=0$, the Hamiltonian (Entanglement patterns of quantum chaotic Hamiltonians with a scalar U(1) charge) becomes the integrable XXZ chain. A finite value of $J_{2}$, instead, breaks integrability. We first consider the parameters $J=1$, $\Delta=J/2$, $J_{2}=0.6$, and $h_{b}=0.25$ which we find to be the values in which eigenstates are most random.